OpenStudy (anonymous):

simplify (x^2-9 over x-5) over x+3

OpenStudy (anonymous):

the equation tool will make this easier to read

OpenStudy (mathstudent55):

Is this the question? \(\large \dfrac{\frac{x^2-9}{x-5}}{x+3} \)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

In general: \(\Large \dfrac{\frac{a}{b} }{\frac{c}{d} }= \dfrac{ab}{bc} \)

OpenStudy (anonymous):

don't simplify divide

OpenStudy (mathstudent55):

Rewrite the denominator of the main fraction as a fraction. All you need to do is write the quantity x + 3 over 1. Now you have a complex fraction of a fraction over a fraction.

OpenStudy (mathstudent55):

Look at the pattern I wrote above. To simplify, multiply the upper most numerator by the lowest denominator and write that in the new numerator. Then multiply the middle denominator and numerator and that goes in the new denominator. Then factor and reduce.

OpenStudy (mathstudent55):